OpenStudy (anonymous):

ABC is similar to PQR. corresponds to , and corresponds to . If AB = 9, BC = 12, CA = 6, and PQ = 3, what are the lengths of and ?

OpenStudy (mathmate):

In geometry, a diagram summarizes and reveals a lot of information:  To find x and y, use equal ratio of corresponding sides: \(\Large \frac{3}{9}=\frac{x}{12}=\frac{y}{6}\)